# On the third homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$

### v1

@@ 2. Low dimensional homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$
$0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0$
@@
By [22, Proposition 4.3] or [24, Theorem 1.2] we have
@@
[…]
it follows that the bottom map is trivial. Note that by the exact sequence obtained just before this proposition, the upper map should be trivial. (i) Let $m$ be even. Then from the above calculation of $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$, we see that $\mu_{2}(A_{m})\otimes H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})=0$ and the natural map $\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ is trivial. Thus (i) follows from Theorem 2.1. (ii) Now assume that $m$ is odd. Then $\mu_{2}(A_{m})\otimes H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\simeq\mathbb{Z}/2$. Moreover, from the commutative diagram
[…]
it follows that the bottom map is surjective (note that the upper map is surjective because $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=0$). Since
@@
for some $\alpha$ and $\beta$ (see [31, Corollary 4.4]), from the group structure of $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ one sees that the natural map $\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ should be injective. Hence the claim follows from Theorem 2.1. ∎
Example 3.2.
Let $A_{m}:=\mathbb{Z}[\frac{1}{m}]$, where $m=2$ or $3$.
(i) The product map $\mu_{2}(A_{2})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})$ is trivial: By [23, Corolllary 3.7, Remark 3.8] we have the isomorphism $H_{3}(\mathit{{\rm PSL}}_{2}(A_{2}),\mathbb{Z})\simeq\mathcal{RB}(A_{2})$, where $\mathcal{RB}(A_{2})$ is the refined Bloch group of $A_{2}$ (see the next section). By a result of Coronado and Hutchinson, we have the refined Bloch-Wigner exact sequence
(see [8, Proposition 8.31]). From these two results it follows that the desired product map is trivial.
(ii) Now we show that the product map $\mu_{2}(A_{3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})$ is not trivial: We know that $H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/12\oplus\mathbb{Z}/4$. This follows from the cohomology calculation of $\mathit{{\rm SL}}_{2}(A_{3})$ in [1, page 8] and Lemma 3.3 below. Furthermore, by [31, Proposition 4.2], $H^{k}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ is generated by $2$-torsion and $3$-torsion elements for $k\geq 3$. Again by Lemma 3.3 it follows that $H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ is generated by $2$-torsion and $3$-torsion elements. Now let the above product map is trivial. Then from Proposition 3.1 we obtain the exact sequence
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0.\end{array}$
This and the isomorphism $H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/12\oplus\mathbb{Z}/4$ imply that $H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ has an element of order $4$, which is a contradiction.
@@
4. A refined Bloch-Wigner exact sequence
Let $A$ be a commutative ring. For any non-negative integer $n$, let $X_{n}(A^{2})$ be the free abelian group generated by the set of all $(n+1)$-tuples $(\langle{\boldsymbol{v}_{0}}\rangle,\dots,\langle{\boldsymbol{v}_{n}}\rangle)$, where every ${\boldsymbol{v}_{i}}\in A^{2}$ is unimodular and for any two distinct vectors ${\boldsymbol{v}_{i}},{\boldsymbol{v}_{j}}$, the matrix $({\boldsymbol{v}_{i}},{\boldsymbol{v}_{j}})$ is invertible. We consider $X_{l}(A^{2})$ as a left $\mathit{{\rm SL}}_{2}(A)$-module (resp. left $\mathit{{\rm PSL}}_{2}(A)$-module) in a natural way. We convert this action to a right action by the definition $m.g:=g^{-1}m$. The $l$-th differential operator
$\partial_{l}:X_{l}(A^{2})\rightarrow X_{l-1}(A^{2}),\ \ l\geq 1,$
is defined as an alternating sum of face operators which throws away the $i$-th component of generators. Let $\partial_{-1}=\epsilon:X_{0}(A^{2})\rightarrow\mathbb{Z}$ be defined by $\sum_{i}n_{i}(\langle{\boldsymbol{v}_{0,i}}\rangle)\mapsto\sum_{i}n_{i}$. Hence we have the complex
$X_{\bullet}(A^{2})\rightarrow\mathbb{Z}:\ \cdots\longrightarrow X_{2}(A^{2})\overset{\partial_{2}}{\longrightarrow}X_{1}(A^{2})\overset{\partial_{1}}{\longrightarrow}X_{0}(A^{2})\overset{\epsilon}{\longrightarrow}\mathbb{Z}\rightarrow 0.$
Proposition 4.1 (Hutchinson).
If $A$ is a local ring, then $X_{\bullet}(A^{2})\rightarrow\mathbb{Z}$ is exact in dimension $<|A/\mathfrak{m}_{A}|$.
Proof.
See [16, Lemma 3.21]. ∎
From now on we assume that $A$ is a local ring. Let $Z_{2}(A^{2}):=\ker(\partial_{2})$ and set
$\mathcal{RP}(A):=H_{0}(\mathit{{\rm SL}}_{2}(A),Z_{2}(A^{2}))=H_{0}(\mathit{{\rm PSL}}_{2}(A),Z_{2}(A^{2})).$
If the residue field of $A$ has at least 4 elements, then $\mathcal{RP}(A)$ is isomorphic with the quotient of the free $\mathcal{G}_{A}$-module generated by the symbols $[x]$, $x\in\mathcal{W}_{A};=\{a\in A:a(1-a)\in{A^{\times}}\}$, over the subgroup generated by the elements
$[x]-[y]+\langle x\rangle\bigg{[}\frac{y}{x}\bigg{]}-\langle x^{-1}-1\rangle\Bigg{[}\frac{1-x^{-1}}{1-y^{-1}}\Bigg{]}+\langle 1-x\rangle\Bigg{[}\frac{1-x}{1-y}\Bigg{]},\ \ \ \ \ \ x,y,x/y\in\mathcal{W}_{A}$
(see [8, §6] and [21, §2]). The inclusion ${\rm inc}:Z_{2}(A^{2})\rightarrow X_{2}(A^{2})$ induces the map
$\lambda:\mathcal{RP}(A)=Z_{2}(A^{2})_{\mathit{{\rm SL}}_{2}(A)}\overset{\overline{{\rm inc}}}{\longrightarrow}X_{2}(A^{2})_{\mathit{{\rm SL}}_{2}(A)}.$
The orbits of the action of $\mathit{{\rm SL}}_{2}(A)$ on $X_{2}(A)$ is represented by
$\langle a\rangle[\ ]:=(\langle{\boldsymbol{e}_{1}}\rangle,\langle{\boldsymbol{e}_{2}}\rangle,\langle{\boldsymbol{e}_{1}}+a{\boldsymbol{e}_{2}}\rangle),\ \ \ \ \langle a\rangle\in\mathcal{G}_{A}.$
Therefore $X_{2}(A^{2})_{\mathit{{\rm SL}}_{2}(A)}\simeq\mathbb{Z}[\mathcal{G}_{A}]$. The group
$\mathcal{RP}_{1}(A):=\ker\Big{(}\lambda:\mathcal{RP}(A)\rightarrow\mathbb{Z}[\mathcal{G}_{A}]\Big{)}.$
is called the scissors congruence group of $A$. We call
${\rm GW}(A):=H_{0}(\mathit{{\rm SL}}_{2}(A),Z_{1}(A^{2}))$
the Grothendieck-Witt group of $A$ (see [21, Remark 2.2] for an explanation on the choice of this name). Let
$\epsilon:=d_{2,0}^{1}:{\rm GW}(A)\rightarrow\mathbb{Z}.$
The kernel of $\epsilon$ is called the fundamental ideal of $A$ and is denoted by $I(A)$. Let $\mathcal{I}_{A}$ be the augmentation ideal of $\mathbb{Z}[\mathcal{G}_{A}]$. It is easy to see that $\mathcal{I}_{A}$ is generated by the elements $\langle\!\langle a\rangle\!\rangle:=\langle a\rangle-1$, $a\in{A^{\times}}$ and the natural map $\mathbb{Z}[\mathcal{G}_{A}]\rightarrow{\rm GW}(A)$ is surjective and maps $\mathcal{I}_{A}$ onto $I(A)$. If $|A/\mathfrak{m}_{A}|\neq 2$, then
${\rm GW}(A)\simeq\mathbb{Z}[\mathcal{G}_{A}]/\langle\langle\!\langle a\rangle\!\rangle\langle\!\langle 1-a\rangle\!\rangle:a,1-a\in{A^{\times}}\rangle,$
$I(A)\simeq\mathcal{I}_{A}/\langle\langle\!\langle a\rangle\!\rangle\langle\!\langle 1-a\rangle\!\rangle:a,1-a\in{A^{\times}}\rangle$
(see [21, §2]). Let $L_{\bullet}(A^{2})$ be the complex $0\rightarrow Z_{1}(A^{2})\overset{{\rm inc}}{\rightarrow}X_{1}(A^{2})\overset{\partial_{1}}{\rightarrow}X_{0}(A^{2})\rightarrow 0$. From this we obtain the first quadrant spectral sequences
$E^{1}_{p.q}=H_{q}(\mathit{{\rm PSL}}_{2}(A),L_{p}(A^{2}))\Longrightarrow H_{p+q}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z}),$
${E^{\prime}}^{1}_{p.q}=H_{q}(\mathit{{\rm SL}}_{2}(A),L_{p}(A^{2}))\Longrightarrow H_{p+q}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}).$
These spectral sequences have been studies extensively in [23] and [21]. One can show that
$E_{0,q}^{1}\simeq H_{q}(\mathit{{\rm PB}}(A),\mathbb{Z}),\ \ \ \ \ \ E_{1,q}^{1}\simeq H_{q}(\mathit{{\rm PT}}(A),\mathbb{Z}),$
${E^{\prime}}_{0,q}^{1}\simeq H_{q}(B(A),\mathbb{Z}),\ \ \ \ \ \ {E^{\prime}}_{1,q}^{1}\simeq H_{q}(T(A),\mathbb{Z}),$
where
$T(A):=\bigg{\{}\begin{pmatrix}a&0\\ 0&a^{-1}\end{pmatrix}:a\in{A^{\times}}\bigg{\}}\ \ \subseteq\ \ B(A):=\bigg{\{}\begin{pmatrix}a&b\\ 0&a^{-1}\end{pmatrix}:a\in{A^{\times}},b\in A\bigg{\}}.$
and
$\mathit{{\rm PB}}(A):=B(A)/\mu_{2}(A)I_{2},\ \ \ \ \ \ \mathit{{\rm PT}}(A):=T(A)/\mu_{2}(A)I_{2}.$
We denote the matrix $\begin{pmatrix}a&0\\ 0&a^{-1}\end{pmatrix}$ by $D(a)$.
Proposition 4.2 (Hutchinson).
Let $A$ be a local ring (local domain) such that either $A/\mathfrak{m}_{A}$ is infinite or if $|A/\mathfrak{m}_{A}|=p^{d}$, we have $(p-1)d>2(n+1)$ ($(p-1)d>2n$). Then $H_{n}(T(A),\mathbb{Z})\simeq H_{n}(B(A),\mathbb{Z})$ and $H_{n}(\mathit{{\rm PT}}(A),\mathbb{Z})\simeq H_{n}(\mathit{{\rm PB}}(A),\mathbb{Z})$.
Proof.
See [21, Proposition 3.8] and [23, Lemma 2.8]. ∎
From now on we assume that $A$ is a local ring (local domain) such that $A/\mathfrak{m}_{A}$ either is infinite or if it has $p^{d}$ elements, then $(p-1)d>8$ ($(p-1)d>6$). Then we have the isomorphisms
$H_{n}(T(A),\mathbb{Z})\simeq H_{n}(B(A),\mathbb{Z})$
$H_{n}(\mathit{{\rm PT}}(A),\mathbb{Z})\simeq H_{n}(\mathit{{\rm PB}}(A),\mathbb{Z})$
for $n\leq 3$. It is shown in [23, Lemma 2.1] that $E_{2,1}^{2}\simeq\mathcal{RP}_{1}(A)$ and $E_{0,2}^{2}\simeq\displaystyle\frac{{A^{\times}}\wedge{A^{\times}}}{{A^{\times}}\wedge\mu_{2}(A)}$. The kernel of
$\lambda_{1}:=d_{2,1}^{2}:\mathcal{RP}_{1}(A)\rightarrow\frac{{A^{\times}}\wedge{A^{\times}}}{{A^{\times}}\wedge\mu_{2}(A)}$
is denoted by $\mathcal{RB}(A)$ and is called the refined Bloch group of $A$ (see [23, Lemma 2.1 and page 8]). Moreover, it is shown in [21, page 11] that the sequence
[…]
is exact. By studying the spectral sequence $E_{\bullet,\bullet}^{1}$ we proved the following projective refined Bloch-Wigner exact sequence in [23].
Theorem 4.3.
Let $A$ be a local ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu(A)\rightarrow\mu(F)$ is injective. Assume that $A/\mathfrak{m}_{A}$ either is infinite or if it has $p^{d}$ elements, then $(p-1)d>8$ (normal-((($(p-1)d>6$ when $A$ is a domain)normal-))). If either $-1\in{A^{\times}}^{2}$ or $|\mathcal{G}_{A}|\leq 4$, then we have the exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(A),\widetilde{\mu}(A))\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mathcal{RB}(A)\rightarrow 0,$
where $\widetilde{\mu}(A)=\mu(A)/\mu_{2}(A)$.
Proof.
See [23, Theorem 3.1, Theorem 3.2, Corollary 3.3] and [21, Proposition 3.8]. ∎
@@ 4. A refined Bloch-Wigner exact sequence
Now we have the following generalization of Hutchinson’s refined Bloch-Wigner exact sequence.
Theorem 4.4.
Let $A$ be a local ring such that there is a ring homomorphism $A\rightarrow F$, $F$ a field, where $\mu(A)\rightarrow\mu(F)$ is injective. Assume that $A/\mathfrak{m}_{A}$ either is infinite or if it has $p^{d}$ elements, then $(p-1)d>8$ (normal-((($(p-1)d>6$ when $A$ is a domain)normal-))). If either $-1\in{A^{\times}}^{2}$ or $|\mathcal{G}_{A}|\leq 4$, then we have the exact sequence
$0\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}\rightarrow\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)}\rightarrow\mathcal{RB}(A)\rightarrow 0$
Proof.
The composite
$\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}\twoheadrightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\twoheadrightarrow\mathcal{RB}(A)$
is surjective. Let $\mathcal{K}$ be the kernel of this map. By Theorem 4.3 we have the commutative diagram with exact rows
[…]
From this diagram and Corollary 2.4 we obtain the exact sequence
$0\rightarrow\mu_{2}(A)^{\sim}\rightarrow\mathcal{K}\rightarrow{{\rm Tor}_{1}^{\mathbb{Z}}}(\widetilde{\mu}(A),\widetilde{\mu}(A))\rightarrow 0$
This clearly implies that $\mathcal{K}\simeq{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu(A),\mu(A))^{\sim}$. Now we study the product map $\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$. Since $A/\mathfrak{m}_{A}$ has at least $4$ elements, we have $A_{A^{\times}}=H_{0}({A^{\times}},A)=0$, where the action of ${A^{\times}}$ on $A$ is given by $a.x:=a^{2}x$. First note that ${E^{\prime}}_{0,1}^{1}\simeq\mathcal{G}_{A}\oplus A_{A^{\times}}=\mathcal{G}_{A}$. The map ${d^{\prime}}_{2,0}^{2}:I(A)\rightarrow\mathcal{G}_{A}$ is given by $\langle\!\langle a\rangle\!\rangle\mapsto\langle a\rangle$ [8, Lemma 4.7]. Thus ${E^{\prime}}_{2,0}^{\infty}\simeq{E^{\prime}}_{2,0}^{2}\simeq I^{2}(A)$. Now from an easy analysis of the spectral sequence ${E^{\prime}}_{\bullet,\bullet}^{1}$ we obtain the exact sequence
$H_{2}(B(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow I^{2}(A)\rightarrow 0.$
By Proposition 4.2, $H_{2}(T(A),\mathbb{Z})\simeq H_{2}(B(A),\mathbb{Z})$. The composite
$\mu_{2}(A)\otimes(T(A)\wedge T(A))\simeq\mu_{2}(A)\otimes H_{2}(T(A),\mathbb{Z})\rightarrow\mu_{2}(A)\otimes H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
$\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})$
is given by $(-1)\otimes(D(a)\wedge D(b))\mapsto{\boldsymbol{c}}(D(-1),D(a),D(b)).$ If $i^{2}=-1$, then
${\boldsymbol{c}}(D(-1),D(a),D(b)=2{\boldsymbol{c}}(D(i),D(a),D(b))=0.$
Let $|\mathcal{G}_{A}|\leq 4$. If $-1\in{A^{\times}}^{2}$, we proceed as in above. If $-1\notin{A^{\times}}^{2}$, then $\mathcal{G}_{A}=\{1,-1,u,-u\}$ for some $u\in{A^{\times}}$. Thus
${\boldsymbol{c}}(D(-1),D(a),D(b))={\boldsymbol{c}}(D(-1),D(-1),D(u))=0.$
In any case, the above composite is trivial and thus
$\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}=\frac{H_{3}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})}{\mu_{2}(A)\otimes_{\mathbb{Z}}I^{2}(A)}.$
This completes the proof of the theorem. ∎

### v3

@@ 3. Low dimensional homology of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$
(4.1) $0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0,$
which will be used in the proof of the following result.
@@
By [19, Proposition 4.3] or [21, Theorem 1.2] we have
@@
${{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(\mathbb{Z}),H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z}))}$${H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})}$${{{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(A_{m}),H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z}))}$${\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}{\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}}$$\scriptstyle{\rho_{\ast}=0}$$\scriptstyle{\rho_{\ast}}$
it follows that the bottom map is trivial. In the above diagram, first note that since $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=0$, we have $H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=\displaystyle\frac{H_{3}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})}{\mu_{2}(\mathbb{Z})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})}$ and thus the above diagram is commutative. Second, by the exact sequence (4.1), the upper map in the above diagram should be trivial.
(i) Let $m$ be even. Then from the above calculation of $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$, we see that
$\mu_{2}(A_{m})\otimes H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})=0$
and the natural map $\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ is trivial. Thus the claim follows from Theorem 3.1.
(ii) Now assume that $m$ is odd. Then $\mu_{2}(A_{m})\otimes H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})\simeq\mathbb{Z}/2$. Moreover, from the commutative diagram
${H_{3}(\mathit{{\rm PSL}}_{2}(\mathbb{Z}),\mathbb{Z})}$${\mu_{2}(\mathbb{Z})\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})}$${H_{3}(\mathit{{\rm PSL}}_{2}(A_{m}),\mathbb{Z})}$${\mu_{2}(A_{m})\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})}$
it follows that the bottom map is surjective (note that the upper map is surjective in view of Theorem 3.1, because $H_{2}(\mathit{{\rm SL}}_{2}(\mathbb{Z}),\mathbb{Z})=0$). Since
@@
for some $\alpha$ and $\beta$ (see [28, Corollary 4.4]), from the group structure of $H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$ one sees that the natural map
$\mu_{2}(A_{m})\rightarrow H_{1}(\mathit{{\rm SL}}_{2}(A_{m}),\mathbb{Z})$
must be injective. Hence the claim follows from Theorem 3.1. ∎
Remark 4.2.
is trivial [13, page 103]. The proof of Theorem 3.1 shows that if $A$ is a ring such that $\mu_{2}(A)=\{\pm 1\}$, then
always is a weak stem-extension. But by Proposition 4.1 this extension is not a stem-extension for $A=\mathbb{Z}[\frac{1}{m}]$, where $m$ is a square free odd integer. This also is true for $A=\mathbb{Z}$ by the paragraph above Proposition 4.1 and $A=\mathbb{Z}_{(2)}$ by Example 5.3.
@@
Example 4.4.
Let $A_{m}:=\mathbb{Z}[\frac{1}{m}]$, where $m=2$ or $3$.
(i) The product map $\mu_{2}(A_{2})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{2}),\mathbb{Z})$ is trivial: By [20, Corolllary 3.7, Remark 3.8] we have the isomorphism
$H_{3}(\mathit{{\rm PSL}}_{2}(A_{2}),\mathbb{Z})\simeq\mathcal{RB}(A_{2}),$
@@
where $\mathcal{RB}(A_{2})$ is the refined Bloch group of $A_{2}$ (see [8, §4] for the definition of $\mathcal{RB}(A_{2})$). By a result of Coronado and Hutchinson, we have the refined Bloch-Wigner exact sequence
(see [8, Proposition 8.31]). From these two results and Proposition 4.1, it follows that the desired product map is trivial.
(ii) Now we show that the product map $\mu_{2}(A_{3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})$ is not trivial. We know that
$H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/12\oplus\mathbb{Z}/4.$
This follows from the cohomology calculation of $\mathit{{\rm SL}}_{2}(A_{3})$ in [1, page 8] and Lemma 4.3. Furthermore, by [28, Proposition 4.2], $H^{k}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ is generated by elements of orders $2$ and $3$ for $k\geq 3$. Again by Lemma 4.3 it follows that $H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ is generated by elements of orders $2$ and $3$. Now let the above product map be trivial. Then by Proposition 4.1 the sequence
$\begin{array}[]{c}0\rightarrow\mathbb{Z}/4\rightarrow H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\rightarrow H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})\rightarrow\mathbb{Z}/2\rightarrow 0\end{array}$
must be exact. This and the isomorphism $H_{3}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/12\oplus\mathbb{Z}/4$ imply that $H_{3}(\mathit{{\rm PSL}}_{2}(A_{3}),\mathbb{Z})$ has an element of order $4$, which is a contradiction. Observe that $H_{2}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}$ and thus $\mu_{2}(A_{3})\otimes_{\mathbb{Z}}H_{2}(\mathit{{\rm SL}}_{2}(A_{3}),\mathbb{Z})\simeq\mathbb{Z}/2$.
5. The low dimensional homology groups of $\mathit{{\rm SL}}_{2}$ and $\mathit{{\rm PSL}}_{2}$ of local rings
Let $A$ be a local ring with maximal ideal $\mathfrak{m}_{A}$. It is known that
$H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\simeq\begin{cases}A/\mathfrak{m}_{A}^{2}&\text{if $|A/\mathfrak{m}_{A}|=2$}\\ A/\mathfrak{m}_{A}&\text{if $|A/\mathfrak{m}_{A}|=3$.}\\ 0&\text{if $|A/\mathfrak{m}_{A}|\geq 4$}\\ \end{cases}$
(see [19, Proposition 4.1]). Now let $|A/\mathfrak{m}_{A}|\neq 2$. Since $\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})=0$, from Theorem 1.1 we obtain the exact sequence
$0\rightarrow H_{2}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})\rightarrow H_{2}(\mathit{{\rm PSL}}_{2}(A),\mathbb{Z})\rightarrow\mu_{2}(A)\rightarrow 0$
Since
$\mu_{2}(A)\otimes_{\mathbb{Z}}H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z})=0,\ \ \ \ {{\rm Tor}_{1}^{\mathbb{Z}}}(\mu_{2}(A),H_{1}(\mathit{{\rm SL}}_{2}(A),\mathbb{Z}))=0,$
by Theorem 3.1 we obtain the following result.
Proposition 5.1.
is exact.